### v1

# Conserved charges and asymptotic symmetries of BTZ-like black holes in Einstein-bumblebee gravity

Abstract
Abstract
Using the solution phase space method, we investigate the conserved charges of BTZ-like black holes in Einstein-bumblebee gravity. Our study shows that the black hole mass, angular momentum and entropy are influenced by the Lorentz-violating parameter. Through the study of the AdS/CFT correspondence, we derive the asymptotic charge algebra, which consists of two copies of the Virasoro algebra with non-trivial central charges. By employing the Cardy formula, we calculate the microscopic entropy of dual conformal field theory, which precisely matches with the Bekenstein-Hawking entropy. Furthermore, by imposing the Detournay-Smoes-Wutte boundary conditions in the near horizon geometry of the extremal BTZ-like black hole, we obtain a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, representing the symmetry algebra of a warped conformal field theory.
Keywords: Conserved charges, Asymptotic symmetries, black hole, Lorentz-violation.
I Introduction
Symmetry is fundamental in modern physics. According to the well known Noether’s theorem, symmetries relate to conservation laws Noether:1918zz . In the realm of black hole physics, the analysis of the asymptotic symmetry of spacetime geometry under appropriate boundary conditions brings new comprehension and great importance. The investigation of asymptotic symmetries in gravitational theories started from the works of Bondi-Metzner-Sachs (BMS) in 1962 Bondi:1962px ; Sachs:1962wk , which identified the BMS group of supertranslations and Lorentz transformations as asymptotic symmetry group (ASG). Another landmark discovery by Brown and Henneaux Brown:1986nw established a definitive link between the asymptotic symmetries of three-dimensional asymptotic anti-de-Sitter ($\mathrm{AdS_{3}}$) spacetime and the symmetries of two-dimensional conformal field theory (CFT), serving as an early precursor of the AdS/CFT correspondence Maldacena:1998bw . By applying the AdS/CFT to the Banados-Teitelboim-Zanelli (BTZ) black holes Banados:1992wn ; Banados:1992gq , the microscopic origin of the Bekenstein-Hawking entropy was derived Strominger:1997eq . A parallel generalization of the AdS/CFT correspondence in near horizon region is the Kerr/CFT correspondence Guica:2008mu , which links near horizon geometry of a four-dimensional extremal Kerr black hole with a two-dimensional chiral CFT.
In the theories of gravity, conserved charges such as black hole mass, angular momentum, electric charge and black hole entropy associated with exact and asymptotic symmetries have been extensively studied by various methods (see Compere:2018aar ; Adami:2017phg ; Frodden:2019ylc for recent reviews of the methods for conserved charges). Based on the Noether procedure, the covariant phase space method (CPSM) was proposed by Wald and Iyer (also called Wald’s formalism) Lee:1990nz ; Wald:1993nt ; Iyer:1994ys . In the framework of CPSM, a systematic approach for calculating conserved charges in generally covariant theories of gravity was developed by K. Hajian et al, which is called solution phase space method (SPSM) Hajian:2015xlp . The SPSM is a restriction of CPSM to solution space manifold, focusing on parametric variation and exact symmetries. In SPSM, the charge variations are computed by an integration over an arbitrary smooth, closed, and compact codimension-2 surface surrounding the black hole singularity. For reviews and applications of SPSM see Refs.Hajian:2016kxx ; Ghodrati:2016vvf ; Ding:2020bwa ; Guo:2024zaa . The study of conserved charges and asymptotic symmetries is crucial for understanding the thermodynamic properties and holographic dualities in gravitational theories. Generally, a physical system is governed by its action and the boundary conditions; the former determines the equations of motion (and therefore the dynamics of the system). This also implies that different actions and boundary conditions contain different conserved charges and ASG.
As a special family of modified gravitational theories, the Lorentz-violating (LV) theories of gravity Mattingly:2005re , including Hořava-Lifshitz theory Horava:2009uw , ghost condensation Arkani-Hamed:2003pdi , warped brane world Frey:2003jq ; Cline:2003xy , Einstein-aether theory Jacobson:2000xp ; Eling:2004dk ; Jacobson:2007veq and Einstein bumblebee gravity Kostelecky:1988zi ; Kostelecky:1989jw ; Kostelecky:2003fs ; Kostelecky:2010ze , have attracted increasing interests. In the action of the bumblebee gravity, the spontaneous Lorentz violation arises from a potential $V(B_{\mu}B^{\mu})$ acting on a bumblebee vector field $B_{\mu}$ Kostelecky:2003fs ; Kostelecky:2010ze . In 2017, Casana et al obtained an exact Schwarzschild-like black hole solution in bumblebee gravity Casana:2017jkc . Subsequently, a number of exact solutions are obtained, including Schwarzschild-anti-de-Sitter (AdS)-like black holes Maluf:2020kgf , slowly rotating Kerr-like black holes Ding:2019mal , rotating BTZ-like black holes Ding:2023niy , and other solutions Santos:2014nxm ; Jha:2020pvk ; Filho:2022yrk ; Xu:2022frb ; Liu:2024axg within bumblebee gravity, all of which exhibit effects induced by LV . Therefore, whether the LV effects in the bumblebee gravity influence the conserved charges and ASG remains an open question worthy of further study.
Although the Lorentz symmetry is broken in the theories of bumblebee gravity, it remains generally invariant under diffeomorphism transformation Bluhm:2007bd . Therefore, the SPSM is still accommodation. In this paper, we employ the SPSM to investigate the conserved charges of a rotating BTZ-like black hole Ding:2023niy in Einstein-bumblebee gravity. We compute the black hole mass, angular momentum and entropy, showing that these quantities are affected by LV parameter. As a consistent check, we also derive the first law of black hole thermodynamics. Mostly, the ASG analysis is carried out in pure gravity Compere:2008cv ; Compere:2009zj ; Blagojevic:2009ek ; Detournay:2023zni ; Godet:2020xpk . In this work, we investigate the ASG in the gravity coupled with the bumblebee field non-minimally. Given that the black hole solution describes an asymptotic AdS spacetime, we study the AdS/CFT correspondence and obtain the asymptotic charge algebra, which consists of two copies of the Virasoro algebra with non-trivial central charges. Using the Cardy formula Cardy:1986ie , we compute the microscopic entropy of dual CFT and find that it is in precise agreement with the Bekenstein-Hawking entropy. Recently, a new class of boundary conditions was devised by Detournay-Smoes-Wutte to study asymptotic symmetries in near horizon geometry Detournay:2023zni , which is a higher dimensional uplift of Godet-Marteau boundary conditions in two-dimensional gravity Godet:2020xpk . An application of these new boundary conditions is given in Ref.Jiang:2024tmh . In near horizon region, we apply the Detournay-Smoes-Wutte boundary conditions to study the asymptotic symmetries for near horizon geometry of extremal BTZ-like black hole. As a result, we obtain a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, which is the symmetry algebra of a warped conformal field theory (WCFT).
This paper is organized as follows. In Sec.2, we provide a brief review of the SPSM. In Sec.3, we derive the Noether-Wald charge density and the surface charge density for Einstein-bumblebee gravity. In Sec.4, we calculate the conserved charges and check the first law of thermodynamics for rotating BTZ-like black hole. In Sec.5, we study the asymptotic symmetries in asymptotic $\mathrm{AdS_{3}}$ background. Two copies of Virasoro charge algebra will be obtained. Subsequently, we derive the microscopic entropy of dual CFT. In the near horizon geometry of extremal BTZ-like black hole, we apply the new boundary conditions to obtain the symmetry algebra of WCFT. Finally, Sec.6 concludes with a summary of our main findings and remarks.
II Solution phase space method
The SPSM Hajian:2015xlp is based on the CPSM Lee:1990nz ; Wald:1993nt ; Iyer:1994ys , so firstly we give a brief review of CPSM to achieve the conserved charges for generic gauge theories.
Covariant phase space method: A phase space is a manifold $\mathcal{M}$ equipped with a symplectic $2$-form $\Omega$. We consider an $n$-dimensional generally covariant theory of gravity described by the action
$S(\Phi)=\int\mathbf{L}(\Phi),$ (1)
where $\mathbf{L}$ is the Lagrangian $n$-form, and $\Phi$ denotes collectively all the dynamical fields. The variation of $\mathbf{L}$ leads to
$\delta\mathbf{L}(\Phi)=\mathbf{E}_{\Phi}\delta\Phi+\mathrm{d}\mathbf{\Theta}(\delta\Phi,\Phi),$ (2)
where $\delta\Phi$ is a generic field variation and provides the basis for tangent bundle of the phase space $\mathcal{M}$. The set $\mathbf{E}_{\Phi}=0$ gives the equations of motion (EOM). $\mathbf{\Theta}$ is an $(n-1;1)$-form called Lee-Wald symplectic potential, where the $(n-1)$-form is over the spacetime and the $1$-form is on the tangent bundle of $\mathcal{M}$. The Lee-Wald symplectic form Lee:1990nz is defined as
$\Omega\left(\delta_{1}\Phi,\delta_{2}\Phi,\Phi\right)=\int_{\Sigma}\bm{\omega}\left(\delta_{1}\Phi,\delta_{2}\Phi,\Phi\right),$ (3)
where
$\bm{\omega}\left(\delta_{1}\Phi,\delta_{2}\Phi,\Phi\right)\equiv\delta_{1}\mathbf{\Theta}\left(\delta_{2}\Phi,\Phi\right)-\delta_{2}\mathbf{\Theta}\left(\delta_{1}\Phi,\Phi\right)$ (4)
representing the symplectic current form, which is an $(n-1;2)$-form. $\Sigma$ is a smooth codimension-$1$ surface.
When $\Phi$ solves the EOM $\mathbf{E}_{\Phi}=0$ and $\delta\Phi$ solves the linearized EOM $\delta\mathbf{E}_{\Phi}=0$, i.e., on-shell, the conserved condition
$\mathrm{d}\bm{\omega}\left(\delta_{1}\Phi,\delta_{2}\Phi,\Phi\right)\approx 0$ (5)
is satisfied, where $\approx$ denotes the on-shell equality.
When the $U(1)$ gauge field $A$ presents in $\mathbf{L}$, in addition to the diffeomorphism generated by a vector field $\xi$, the Lagrangian remains gauge invariant under the transformation $A\rightarrow A+\mathrm{d}\lambda$ for an arbitrary scalar $\lambda$. Consequently, the general gauge transformation takes the form $\delta_{\epsilon}\Phi=\left\{\mathcal{L}_{\xi}\Phi,\delta_{\lambda}\Phi\right\}$ where the generator $\epsilon$ is a combination of diffeomorphism and gauge transformation, given by $\epsilon\equiv\left\{\xi,\lambda\right\}$.
For field-dependent transformations, i.e. $\delta\epsilon\neq 0$, the charge variation $\delta H_{\epsilon}(\Phi)$ associated with generators $\epsilon$ can be defined as
$\delta H_{\epsilon}(\Phi)$ $\equiv\Omega\left(\delta\Phi,\delta_{\epsilon}\Phi,\Phi\right)=\int_{\Sigma}\bm{\omega}\left(\delta\Phi,\delta_{\epsilon}\Phi,\Phi\right)$
$\equiv\int_{\Sigma}\left(\delta^{[\Phi]}\mathbf{\Theta}\left(\delta_{\epsilon}\Phi,\Phi\right)-\delta_{\epsilon}\mathbf{\Theta}(\delta\Phi,\Phi)\right)$
$=\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\delta\Phi,\Phi),$ (6)
where
$\bm{\omega}\left(\delta\Phi,\delta_{\epsilon}\Phi,\Phi\right)\approx\mathrm{d}\bm{k}_{\epsilon}(\delta\Phi,\Phi),$ (7)
and
$\bm{k}_{\epsilon}(\delta\Phi,\Phi)=\delta\mathbf{Q}_{\epsilon}-\xi\cdot\mathbf{\Theta}(\delta\Phi,\Phi).$ (8)
Here $\bm{k}_{\epsilon}$ is an $(n-2;1)$-form, called the surface charge density. The $(n-2;2)$-form $\mathbf{Q}_{\epsilon}$ is Noether-Wald charge density, related to the Noether current $\mathbf{J}_{\epsilon}$ by $\mathbf{J}_{\epsilon}\approx\mathrm{d}\mathbf{Q}_{\epsilon}$, and
$\mathrm{d}\mathbf{Q}_{\epsilon}\equiv\mathbf{\Theta}\left(\delta_{\epsilon}\Phi,\Phi\right)-\xi\cdot\mathbf{L}.$ (9)
In Eq.(II), $\delta^{[\Phi]}$ denotes that $\delta$ acts only on $\Phi$ but not on the $\epsilon$ inside $\mathbf{\Theta}$.
If $\delta H_{\epsilon}(\Phi)$ is well-defined and integrable, we could determine the conserved charge $H_{\epsilon}$. The integrability condition is basically $\left(\delta_{1}\delta_{2}-\delta_{2}\delta_{1}\right)H_{\epsilon}(\Phi)=0$, which is equivalent to Hajian:2015xlp
$\oint_{\partial\Sigma}\left(\xi\cdot\bm{\omega}\left(\delta_{1}\Phi,\delta_{2}\Phi,\Phi\right)+\bm{k}_{\delta_{1}\epsilon}\left(\delta_{2}\Phi,\Phi\right)-\bm{k}_{\delta_{2}\epsilon}\left(\delta_{1}\Phi,\Phi\right)\right)\approx 0.$ (10)
To ensure the conservation of charge variation $\delta H_{\epsilon}(\Phi)$ and the independence of integration on $\Sigma$ and $\partial\Sigma$, the symplectic current form $\bm{\omega}$ must vanish on-shell for a certain subclass of $\delta_{\epsilon}\Phi$’s, i.e.,
$\bm{\omega}\left(\delta\Phi,\delta_{\epsilon}\Phi,\Phi\right)\approx 0.$ (11)
In this case the transformations generated by $\delta_{\epsilon}\Phi$ are called symplectic symmetries. The family of $\epsilon$’s with this property can be divided into two sets: (1) non-exact symmetry generators denoted by $\chi$, for which $\delta_{\chi}\Phi\neq 0$ at least on one point in the phase space. (2) exact symmetry generators denoted by $\eta$, for which $\delta_{\eta}\Phi=0$ all over the phase space.
Solution phase space method: The SPSM is a specification of CPSM to solution space manifold, which deals with parametric variation and exact symmetries $\hat{\delta}_{\eta}\Phi=0$. The solution space manifold $\hat{\mathcal{M}}$ is composed of solutions $\hat{\Phi}=\hat{\Phi}\left(x;p_{\alpha}\right)$, where $p_{\alpha}$ is a set of solution parameters. The parametric variations $\hat{\delta}\Phi$ are defined as
$\hat{\delta}\Phi\equiv\frac{\partial\hat{\Phi}}{\partial p_{\alpha}}\delta p_{\alpha}.$ (12)
The corresponding symplectic $2$-form $\Omega$ is denoted as $\hat{\Omega}$. The solution phase space is denoted by (^$(\hat{\mathcal{M}},\hat{\Omega})$, which is a submanifold of $(\mathcal{M},\Omega)$ when $\{\Phi\}$ are restricted to solution space $\{\hat{\Phi}\}$. The charge variations associated with the generators $\epsilon$ are taken as
$\hat{\delta}H_{\epsilon}(\Phi)=\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\hat{\delta}\Phi,\hat{\Phi}).$ (13)
The integrability condition (10) for parametric variations is given by
$\oint_{\partial\Sigma}(\xi\cdot\bm{\omega}(\hat{\delta}_{1}\Phi,\hat{\delta}_{2}\Phi,\hat{\Phi})+\bm{k}_{\hat{\delta}_{1}\epsilon}(\hat{\delta}_{2}\Phi,\hat{\Phi})-\bm{k}_{\hat{\delta}_{2}\epsilon}(\hat{\delta}_{1}\Phi,\hat{\Phi}))\approx 0.$ (14)
If the integration of Eq.(13) is well-defined and integrable over parameters $p_{\alpha}$, the conserved charges $H_{\epsilon}\left(p_{\alpha}\right)$ could be calculated by
$H_{\epsilon}(\hat{\Phi}(p_{\alpha}),\bar{\Phi}(\bar{p}_{\alpha}))=\int_{\bar{p}}^{p}\hat{\delta}H_{\epsilon}+H_{\epsilon}(\bar{\Phi}(\bar{p}_{\alpha})),$ (15)
where the integration is performed over arbitrary integral curves connecting a reference (or background) field configuration $\bar{\Phi}(x^{\mu};\bar{p}_{\alpha})$ to the target space $\Phi(x^{\mu};p_{\alpha})$. $H_{\epsilon}(\bar{\Phi}(\bar{p}_{\alpha}))$ is the reference point for $H_{\epsilon}$ defined on $\bar{\Phi}(x^{\mu};\bar{p}_{\alpha})$.
Algebra of surface charges: When the charge variation $\delta H_{\epsilon}(\Phi)$ is well-defined and integrable, the algebra of the asymptotic symmetries is given by the Poisson bracket algebra of the surface charges Barnich:2007bf ; Compere:2007az
$i\left\{H_{\epsilon}(\Phi,\bar{\Phi}),H_{\tilde{\epsilon}}(\Phi,\bar{\Phi})\right\}$ $\equiv i\delta_{\tilde{\epsilon}}H_{\epsilon}(\Phi)$
$=i\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\delta_{\tilde{\epsilon}}\Phi,\Phi)$
$=iH_{[\epsilon,\tilde{\epsilon}]}(\Phi,\bar{\Phi})-iH_{[\epsilon,\tilde{\epsilon}]}(\bar{\Phi})+i\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\delta_{\tilde{\epsilon}}\bar{\Phi},\bar{\Phi}),$ (16)
where $[\epsilon,\tilde{\epsilon}]$ denotes the Lie bracket of the generator $\epsilon$, and the last term $i\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\delta_{\tilde{\epsilon}}\bar{\Phi},\bar{\Phi})$ is the central extention term in the algebra.
III Einstein-bumblebee theory
In the bumblebee theory of gravity, the bumblebee vector field $B_{\mu}$ is introduced to couples with gravity non-minimally. This field acquires a nonzero vacuum expectation value, leading to the spontaneous breaking of Lorentz symmetry in the gravitational sector. For Einstein-bumblebee gravity, the dynamical fields $\Phi$ consist of the metric $g_{\mu\nu}$ and bumblebee field $B_{\mu}$. The Lagrangian is given by Ding:2023niy
$L=\frac{1}{2\kappa}\left(R-2\Lambda\right)+\frac{\rho}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B_{\mu}B^{\mu}\pm b^{2}),$ (17)
where $R$ is the Ricci scalar, and $\Lambda=-1/l^{2}$ is the negative cosmological constant. $\kappa=4\pi G$ is the gravitational coupling constant, with $G$ being the Newtonian constant. $\rho$ is the nonminimal coupling constant between gravity and the bumlebee field. Analogous to the electromagnetic field, the bumblebee field strength is defined as
$B_{\mu\nu}=\nabla_{\mu}B_{\nu}-\nabla_{\nu}B_{\mu}.$ (18)
The potential $V(B_{\mu}B^{\mu}\pm b^{2})$ triggers Lorentz violation, leading to a nonvanishing vacuum expectation value for $B_{\mu}$, such that $B^{\mu}\right>=b^{\mu}$, where the vector $b_{\mu}$ satisfies $b^{\mu}b_{\mu}=\mp b^{2}=const$.
The Lagrangian $n$-form is the Hodge dual of $L$,
$\mathbf{L}=\star{L}=L\,\bm{\epsilon},$ (19)
where $\bm{\epsilon}=\sqrt{-g}\,d^{n}x=\frac{\sqrt{-g}}{n!}\varepsilon_{\mu_{1}\cdots\mu_{n}}dx^{\mu_{1}}\wedge\cdots\wedge dx^{\mu_{n}}$ is the $n$-dimensional volume form, $g$ is the determinant of the metric $g_{\mu\nu}$, and $\varepsilon_{\mu_{1}\cdots\mu_{n}}$ is the Levi-Civita symbol. The variations of the Lagriangian $n$-form $\mathbf{L}$ with respect to $g_{\mu\nu}$ and $B_{\mu}$ are
$\delta\mathbf{L}(\Phi)=\left(\mathrm{E}^{\mu\nu}_{g}\delta g_{\mu\nu}+\mathrm{E}^{\mu}_{B}\delta B_{\mu}\right)\bm{\epsilon}+\mathrm{d}\mathbf{\Theta}(\delta\Phi,\Phi)$ (20)
where the EOMs
$\mathrm{E}^{\mu\nu}_{g}=$ $\frac{1}{2\kappa}\left[-(G^{\mu\nu}+\Lambda g^{\mu\nu})\right]$
$+$ $\frac{\rho}{2\kappa}\bigg{[}\frac{1}{2}g^{\mu\nu}B^{\sigma}B^{\rho}R_{\sigma\rho}-2B^{(\nu}R^{\mu)}{}_{\sigma}B^{\sigma}+\nabla_{\sigma}\nabla^{(\mu}(B^{\nu)}B^{\sigma})$
-- $\frac{1}{2}\nabla_{\sigma}\nabla^{\sigma}(B^{\mu}B^{\nu})-\frac{1}{2}g^{\mu\nu}\nabla_{\sigma}\nabla_{\rho}(B^{\sigma}B^{\rho})\bigg{]}$
$+$ $\frac{1}{2}\left[B^{(\mu}{}_{\sigma}B^{\nu)\sigma}-\frac{1}{4}g^{\mu\nu}B^{\sigma\rho}B_{\sigma\rho}-g^{\mu\nu}V+2V^{\prime}B^{\mu}B^{\nu}\right],$
$\mathrm{E}^{\mu}_{B}=$ $\frac{\rho}{2\kappa}\left[2g^{\mu\nu}B^{\sigma}R_{\nu\sigma}\right]+\nabla_{\nu}B^{\nu\mu}-2V^{\prime}B^{\mu},$ (21)
in which
$G_{\mu\nu}=$ $R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu},$
$V^{\prime}=$ $\frac{\partial V(X)}{\partial X},$
$X=$ $B_{\mu}B^{\mu}\pm b^{2}.$ (22)
The symplectic potential $\mathbf{\Theta}(\delta\Phi,\Phi)$ is given by
$\mathbf{\Theta}(\delta\Phi,\Phi)=\Theta^{\mu}(\delta\Phi,\Phi)\sqrt{-g}(d^{n-1}x)_{\mu},$ (23)
in which
$\Theta^{\mu}(\delta\Phi,\Phi)=\frac{1}{2\kappa}\left[2\nabla^{[\nu}{h^{\mu]}}_{\nu}\right]+\frac{\rho}{2\kappa}\left[\nabla_{\nu}(\Omega^{\mu\nu\sigma\rho}\delta g_{\sigma\rho})+\Sigma^{\mu\sigma\nu}\delta g_{\sigma\nu}\right]-B^{\mu\nu}\delta B_{\nu},$ (24)
with
$\Omega^{\mu\nu\rho\sigma}=$ $g^{\mu(\rho}B^{\sigma)}B^{\nu}-\frac{1}{2}g^{\mu\nu}B^{\sigma}B^{\rho}-\frac{1}{2}g^{\sigma\rho}B^{\mu}B^{\nu},$
$\Sigma^{\mu\sigma\nu}=$ $-g^{\mu(\nu}\nabla_{\rho}(B^{\sigma)}B^{\rho})-\nabla^{(\sigma}(B^{\nu)}B^{\mu})+\nabla^{\mu}(B^{\sigma}B^{\nu})+g^{\nu\sigma}\nabla_{\rho}(B^{\rho}B^{\mu}).$ (25)
Here and in what follows $h_{\mu\nu}=\delta g_{\mu\nu},\,\,\,\,h^{\mu\nu}=g^{\mu\alpha}g^{\nu\beta}\delta g_{\alpha\beta}=-\delta g^{\mu\nu},\,\,\,\,h=g^{\mu\nu}\delta g_{\mu\nu}$, and $\left(d^{n-p}x\right)_{\mu_{1}\cdots\mu_{p}}=\frac{1}{p!(n-p)!}\varepsilon_{\mu_{1}\cdots\mu_{p}\nu_{p+1}\cdots\nu_{n}}dx^{\nu_{p+1}}\wedge\cdots\wedge dx^{\nu_{n}}$.
In the bumblebee gravity, the $\mathrm{U}(1)$ symmetry is broken, thus the transformation is generated by diffeomorphism $\xi$ only, i.e., $\epsilon=\xi$. The fields transformation are
$\delta_{\xi}g_{\mu\nu}$ $=\mathcal{L}_{\xi}g_{\mu\nu}=\nabla_{\mu}\xi_{\nu}+\nabla_{\nu}\xi_{\mu},$ (26)
$\delta_{\xi}B_{\mu}$ $=\mathcal{L}_{\xi}B_{\mu}=\xi^{\nu}\nabla_{\nu}B_{\mu}+B_{\nu}\nabla_{\mu}\xi^{\nu},$
where $\mathcal{L}_{\xi}$ is the Lie derivative along the vector field $\xi$. From Eq.(9) we can get the Noether-Wald charge density
$\mathbf{Q}_{\xi}=\mathrm{Q}^{\mu\nu}_{\xi}\sqrt{-g}(d^{n-2}x)_{\mu\nu},$ (27)
where
$\mathrm{Q}^{\mu\nu}_{\xi}=$ $\frac{1}{2\kappa}\left[-2\nabla^{[\mu}\xi^{\nu]}\right]$
$+$ $\frac{\rho}{2\kappa}\left[2\nabla^{[\mu}(B^{\nu]}B_{\sigma})\xi^{\sigma}+4\nabla_{\sigma}(B^{\sigma}B^{[\mu})\xi^{\nu]}-\nabla_{\sigma}(2B^{\sigma}B^{[\mu}\xi^{\nu]})\right]$
-- $B^{\mu\nu}B_{\sigma}\xi^{\sigma}.$ (28)
Thus, by using Eq.(8) we get the surface charge density
$\bm{k}_{\xi}(\delta\Phi,\Phi)=\sqrt{-g}k^{\mu\nu}_{\xi}(d^{n-2}x)_{\mu\nu}\,,$ (29)
where
$k^{\mu\nu}_{\xi}(\delta\Phi,\Phi)=k^{\mu\nu}_{g}(\delta\Phi,\Phi)+k^{\mu\nu}_{gB}(\delta\Phi,\Phi)+k^{\mu\nu}_{B}(\delta\Phi,\Phi),$ (30)
with
$k^{\mu\nu}_{g}(\delta\Phi,\Phi)=$ $\frac{1}{2\kappa}\cdot 2\left[-\frac{1}{2}h\nabla^{[\mu}\xi^{\nu]}+h^{\sigma[\mu}\nabla_{\sigma}\xi^{\nu]}-\xi_{\sigma}\nabla^{[\mu}h^{\nu]\sigma}-\xi^{[\mu}\nabla^{\nu]}h+\xi^{[\mu}\nabla_{\sigma}h^{\nu]\sigma}\right],$
$k^{\mu\nu}_{gB}(\delta\Phi,\Phi)=$ $\frac{\rho}{2\kappa}\cdot 2\bigg{[}\frac{1}{2}h\nabla^{[\mu}(B^{\nu]}B_{\sigma})\xi^{\sigma}+h\nabla_{\sigma}(B^{\sigma}B^{[\mu})\xi^{\nu]}-\frac{1}{2}h\nabla_{\sigma}(B^{\sigma}B^{[\mu}\xi^{\nu]})-h^{\alpha[\mu}\nabla_{\alpha}(B^{\nu]}B_{\sigma})\xi^{\sigma}$
$+g^{\alpha[\mu}\delta[\nabla_{\alpha}(B^{\nu]}B_{\sigma})]\xi^{\sigma}+2\delta[\nabla_{\sigma}(B^{\sigma}B^{[\mu})]\xi^{\nu]}-\delta[\nabla_{\sigma}(B^{\sigma}B^{[\mu}\xi^{\nu]})]$
$-[\nabla_{\alpha}(\Omega^{[\mu|\alpha\sigma\rho|}h_{\sigma\rho})+\Sigma^{[\mu|\alpha\sigma|}h_{\alpha\sigma}]\cdot\xi^{\nu]}\bigg{]},$
$k^{\mu\nu}_{B}(\delta\Phi,\Phi)=$ $2\left[-\frac{1}{4}hB^{\mu\nu}B_{\sigma}\xi^{\sigma}-\frac{1}{2}\delta B^{\mu\nu}B_{\sigma}\xi^{\sigma}-\frac{1}{2}B^{\mu\nu}\delta B_{\sigma}\xi^{\sigma}-B^{\sigma[\mu}\delta B_{\sigma}\xi^{\nu]}\right],$ (31)
in which
$\delta B^{\mu\nu}=2h_{\alpha}{}^{[\mu}B^{\nu]\alpha}+2g^{\alpha[\mu}g^{\nu]\beta}\nabla_{\alpha}[\delta B_{\beta}],$
$\delta[\nabla_{\alpha}(B^{\nu}B_{\sigma})]=-B_{\sigma}h^{\nu}{}_{\beta}\nabla_{\alpha}B^{\beta}+\delta B_{\sigma}\nabla_{\alpha}B^{\nu}+\delta B^{\nu}\nabla_{\alpha}B_{\sigma}-B^{\beta}h^{\nu}{}_{\beta}\nabla_{\alpha}B_{\sigma}+B_{\sigma}\nabla_{\alpha}\delta B^{\nu}$
$+B^{\nu}\nabla_{\alpha}\delta B_{\sigma}-B^{\nu}B_{\beta}\delta\Gamma^{\beta}{}_{\alpha\sigma}-g^{\nu\rho}B_{\sigma}B_{\beta}\delta\Gamma^{\beta}{}_{\alpha\rho},$
$\delta[\nabla_{\sigma}(B^{\sigma}B^{\mu})]=2B^{(\alpha}\nabla_{\alpha}\delta B^{\mu)}+2\delta B^{(\alpha}\nabla_{\alpha}B^{\mu)}-2B^{\alpha}h^{\mu}{}_{(\alpha}\nabla_{\beta)}B^{\beta}-2h_{\alpha}{}^{\beta}B^{(\alpha}\nabla_{\beta}B^{\mu)}$
$-B_{\alpha}B^{\mu}\nabla_{\beta}h^{\alpha\beta}+\frac{1}{2}B^{\alpha}B^{\mu}\nabla_{\alpha}h-\frac{1}{2}B^{\alpha}B^{\beta}\nabla^{\mu}h_{\alpha\beta},$
$\delta[\nabla_{\sigma}(B^{\sigma}B^{[\mu}\xi^{\nu]})]=-B^{\alpha}h_{\beta}{}^{[\mu}\xi^{\nu]}\nabla_{\alpha}B^{\beta}+\delta B^{\alpha}\xi^{[\nu}\nabla_{\alpha}B^{\mu]}+B^{\alpha}\xi^{[\nu}\nabla_{\alpha}\delta B^{\mu]}+\frac{1}{2}B^{\alpha}B^{[\mu}\xi^{\nu]}\nabla_{\alpha}h$
$+\frac{1}{2}B^{\alpha}\xi^{\beta}B^{[\mu}\nabla_{\alpha}h^{\nu]}{}_{\beta}+\delta B^{\alpha}B^{[\mu}\nabla_{\alpha}\xi^{\nu]}+B^{\alpha}\delta B^{[\mu}\nabla_{\alpha}\xi^{\nu]}-B^{\alpha}B^{\beta}h^{[\mu}{}_{\beta}\nabla_{\alpha}\xi^{\nu]}+\delta B^{[\mu}\xi^{\nu]}\nabla_{\alpha}B^{\alpha}$
$-B^{\alpha}h^{[\mu}{}_{\alpha}\xi^{\nu]}\nabla_{\beta}B^{\beta}-B^{\alpha}B^{[\mu}\xi^{\nu]}\nabla^{\beta}h_{\alpha\beta}+\frac{1}{2}B^{\alpha}\xi^{\beta}B^{[\mu}\nabla_{\beta}h^{\nu]}{}_{\alpha}-h_{\alpha\beta}B^{[\mu}\xi^{\nu]}\nabla^{\beta}B^{\alpha}+B^{[\mu}\xi^{\nu]}\nabla_{\alpha}\delta B^{\alpha}$
$-h_{\alpha}{}^{\beta}B^{\alpha}\xi^{[\nu}\nabla_{\beta}B^{\mu]}-h_{\alpha}{}^{\beta}B^{\alpha}B^{[\mu}\nabla_{\beta}\xi^{\nu]}-\frac{1}{2}B^{\alpha}B^{\beta}\xi^{[\nu}\nabla^{\mu]}h_{\alpha\beta}-\frac{1}{2}B^{\alpha}\xi^{\beta}B^{[\mu}\nabla^{\nu]}h_{\alpha\beta}.$ (32)
IV Conserved charges for BTZ-like bumblebee black hole
When the potential $V=\frac{\lambda}{2}X$ in the Lagrangian, where $\lambda$ is a Lagrange-multiplier, and there is no bumblebee field frosted at its vacuum expectation value, i.e., $B_{\mu}=b_{\mu}$, the $3$-dimensional rotating BTZ-like bumblebee black hole solution is given by Ding:2023niy
$ds^{2}=$ $-f(r)dt^{2}+\frac{(1+s)}{f(r)}dr^{2}+r^{2}(d\phi-\frac{j}{2r^{2}}dt)^{2},$
$f(r)=$ $-\mu-\Lambda r^{2}+\frac{j^{2}}{4r^{2}},$
$B_{\mu}=$ $(0,~{}B_{r},~{}0),$
$B_{r}=$ $b_{0}\left[\frac{(1+s)}{f(r)}\right]^{\frac{1}{2}},$ (33)
where $\mu$ and $j$ are mass and angular momentum parameters, respectively. $s=\rho b^{2}_{0}$ represents the spontaneous Lorentz symmetry breaking parameter, with $b_{0}$ being a positive constant. The negative cosmological constant $\Lambda=-\frac{1}{l^{2}}$ is related to the Lagrange-multiplier $\lambda$, $\Lambda=(1+s)\frac{\kappa\lambda}{2\rho}$.
This solution has two horizons: inner horizon $r_{-}$ and out horizon $r_{+}$, which solve $f(r_{\pm})=0$,
$r_{\pm}=\frac{\sqrt{2}}{2}l\sqrt{\mu\pm\sqrt{\mu^{2}-\frac{j^{2}}{l^{2}}}}.$ (34)
Since the integration in Eq.(13) is independent of the choice of integration surface $\partial\Sigma$, we take $\partial\Sigma$ to be the circle $S$ of constant $(t,r)$ for simplicity and take the limit $r\rightarrow\infty$. Then, the conserved charge variation in $3$-dimensional gravity associated with exact symmetry $\eta$ can be
$\hat{\delta}H_{\eta}=\oint_{S}\bm{k}_{\eta}(\hat{\delta}\Phi,\hat{\Phi})=\int_{0}^{2\pi}\lim_{r\to\infty}\sqrt{-\hat{g}}\,k_{\eta}^{tr}(\hat{\delta}\Phi,\hat{\Phi})\mathrm{d}\phi,$ (35)
where $k_{\eta}^{tr}$ is the $tr$ component of $k_{\eta}^{\mu\nu}$. The dynamical fields are $\hat{\Phi}=(\hat{g}_{\mu\nu},\hat{B}_{\mu})$, parametrized by $p_{\alpha}=\{\mu,j\}$. The parametric variations are given by
$\hat{\delta}g_{\mu\nu}=$ $\frac{\partial\hat{g}_{\mu\nu}}{\partial\mu}\delta\mu+\frac{\partial\hat{g}_{\mu\nu}}{\partial j}\delta j,$
$\hat{\delta}B_{\mu}=$ $\frac{\partial\hat{B}_{\mu}}{\partial\mu}\delta\mu+\frac{\partial\hat{B}_{\mu}}{\partial j}\delta j.$ (36)
The stationary axial symmetric black hole metric has two killing vectors $\xi=\partial_{t}$ and $\xi=\partial_{\phi}$, which are exact symmetries of the spacetime, and any linear combination of these two with $\mu,j$-dependent coefficients is also a killing.
Mass: By choosing the exact symmetry $\eta_{{}_{M}}=\partial_{t}$ and substituting the surface charge expressions (29)-(III) along with the parametric variations (IV) into Eq.(35), we can obtain the total mass
$\hat{\delta}M=\hat{\delta}H_{\eta_{{}_{\mathrm{M}}}}=\frac{\sqrt{1+s}\delta\mu}{4G}\quad\Rightarrow\quad M=\frac{\sqrt{1+s}\mu}{4G}.$ (37)
Angular momentum: By choosing the exact symmetry $\eta_{{}_{J}}=-\partial_{\phi}$, and using Eq.(35), the angular momentum is given by
$\hat{\delta}J=\hat{\delta}H_{\eta_{{}_{\mathrm{J}}}}=\frac{\sqrt{1+s}\delta j}{4G}\quad\Rightarrow\quad J=\frac{\sqrt{1+s}j}{4G}.$ (38)
Entropy: The killing vector generating the killing horizon of the black hole is
$\zeta=\partial_{t}+\Omega_{\mathrm{H}}\partial_{\phi}.$ (39)
The surface gravity, temperature and angular velocity on the horizon are given by Compere:2018aar
$\kappa_{{}_{\mathrm{H}}}=$ $\sqrt{-\frac{1}{2}(\nabla_{\mu}\zeta_{\nu})(\nabla^{\mu}\zeta^{\nu})}|_{r_{{}_{\mathrm{H}}}}=\frac{r^{2}_{\pm}-r^{2}_{\mp}}{\sqrt{1+s}l^{2}r_{\pm}},$
$T_{\mathrm{H}}=$ $\frac{\kappa_{{}_{\mathrm{H}}}}{2\pi},$
$\Omega_{\mathrm{H}}=$ $-\frac{g_{t\varphi}}{g_{\varphi\varphi}}|_{r_{{}_{\mathrm{H}}}}=\frac{j}{2{r^{2}_{{}_{\mathrm{H}}}}}.$ (40)
where $r_{{}_{\mathrm{H}}}\equiv r_{\pm}$. By choosing the exact symmetry $\eta_{{}_{\mathrm{H}}}=\frac{2\pi}{\kappa_{{}_{\mathrm{H}}}}\zeta$, and using Eq.(35), we obtain the entropy
$\hat{\delta}S_{\mathrm{H}}=\hat{\delta}H_{\eta_{{}_{\mathrm{H}}}}=\frac{(1+s)\pi\delta r_{{}_{\mathrm{H}}}}{G}\quad\Rightarrow\quad S_{\mathrm{H}}=\frac{(1+s)\pi r_{{}_{\mathrm{H}}}}{G}.$ (41)
The first law: With the decomposition
$\eta_{{}_{\mathrm{H}}}=\frac{1}{T_{H}}(\eta_{{}_{M}}-\Omega_{\mathrm{H}}\eta_{{}_{J}}),$ (42)
and the linearity of $\hat{\delta}H_{\eta_{{}_{\mathrm{H}}}}$ in $\eta$, the first law is satisfied
$\delta S=\frac{1}{T_{\mathrm{H}}}(\delta M-\Omega_{\mathrm{H}}\delta J).$ (43)
In our formulation the reference points for the charges above are chosen to vanish when $\mu=j=0$. The total mass, angular momentum and entropy are all affected by the factor $\sqrt{1+s}$, arising due to LV effect. The first law is satisfied perfectly. When $s\rightarrow 0$ , this solution reduces to the exact BTZ black hole solution, and $M=\frac{\mu}{4G}$, $J=\frac{j}{4G}$, $S=\frac{\pi r_{{}_{\mathrm{H}}}}{G}$ (up to factor $\frac{1}{2}$ , which comes from the choice of the gravitational coupling constant $\kappa=4\pi G$ in original black hole solution Ding:2023niy and also in our case). As a consistent check, in the next section we will derive the microscopic entropy of dual CFT by the AdS/CFT correspondence, and find that the Bekenstein-Hawking entropy (41) is in precise agreement with the microscopic entropy of the dual CFT.
V Asymptotic symmetries
In this section we will analysis the asymptotic symmetries in $\mathrm{AdS}_{3}$ background, and obtain the warped conformal algebra in the near horizon region by employing the new boundary conditions.
V.1 AdS/CFT correspondence
Setting $\mu=0=j$ in the metric and bumblebee field of the BTZ-like black hole (IV), we can get the $\mathrm{AdS}_{3}$ background spacetime
$d\bar{s}^{2}=$ $-\frac{r^{2}}{l^{2}}dt^{2}+\frac{(1+s)l^{2}}{r^{2}}dr^{2}+r^{2}d\phi^{2},$
$\bar{B}_{\mu}=$ $\left(0,~{}\frac{b_{0}\sqrt{1+s}l}{r},~{}0\right).$ (44)
By using the Brown-Henneaux boundary conditions Strominger:1997eq
$\delta g_{\mu\nu}=\mathcal{O}\left(\begin{array}[]{ccc}1{}{}{}&\frac{1}{r^{3}}{}{}{}&1\\ &\frac{1}{r^{4}}{}{}{}&\frac{1}{r^{3}}\\ &&1\end{array}\right),$ (45)
and choose the following boundary conditions for $B_{\mu}$,
$\delta B_{\mu}=\mathcal{O}\left(1,~{}\frac{1}{r^{3}},~{}1\right),$ (46)
one can get the asymptotic killing vectors $\xi=\xi^{\mu}\partial_{\mu}$, which preserve (45) and (46),
$\xi^{t}=$ $l(\mathrm{T}^{+}+\mathrm{T}^{-})+\frac{l^{3}}{2r^{2}}(\partial^{2}_{+}\mathrm{T}^{+}+\partial^{2}_{-}\mathrm{T}^{-})+\mathcal{O}(\frac{1}{r^{4}})$
$\xi^{\phi}=$ $\mathrm{T}^{+}-\mathrm{T}^{-}-\frac{l^{2}}{2r^{2}}(\partial^{2}_{+}\mathrm{T}^{+}-\partial^{2}_{-}\mathrm{T}^{-})+\mathcal{O}(\frac{1}{r^{4}})$
$\xi^{r}=$ $-r(\partial_{+}\mathrm{T}^{+}+\partial_{-}\mathrm{T}^{-})+\mathcal{O}(\frac{1}{r}),$ (47)
where $x^{\pm}=\frac{t}{l}\pm\phi$, $2\partial_{\pm}=l\partial_{t}\pm\partial_{\phi}$, and $\mathrm{T}^{\pm}=\mathrm{T}^{\pm}(x^{\pm})$. By taking modes $\xi^{\pm}_{m}=\xi(\mathrm{T}^{\pm}=\frac{1}{2}e^{imx^{\pm}})$, $m\in\mathbb{Z}$, one can get the Witt algebras through Lie brackets,
$i\left[\xi^{\pm}_{m},\xi^{\pm}_{n}\right]=$ $(m-n)\xi^{\pm}_{(m+n)},$
$i\left[\xi^{+}_{m},\xi^{-}_{n}\right]=$ $0$ (48)
For the generator $\xi^{+}_{m}$, the Poisson bracket algebra (II) becomes
$i\left\{\tilde{L}^{+}_{m},\tilde{L}^{+}_{n}\right\}=(m-n)\tilde{L}^{+}_{m+n}-iH_{[\xi^{+}_{m},\xi^{+}_{n}]}(\bar{\Phi})+i\oint_{S}\bm{k}_{\xi^{+}_{m}}(\delta_{\xi^{+}_{n}}\bar{\Phi},\bar{\Phi}),$ (49)
where the central extention term
$i\oint_{S}\bm{k}_{\xi^{+}_{m}}(\delta_{\xi^{+}_{n}}\bar{\Phi},\bar{\Phi})=\frac{1}{12}\frac{3\sqrt{1+s}l}{G}m^{3}\delta_{m+n,0}.$ (50)
Similarly, for $i\left\{\tilde{L}^{-}_{m},\tilde{L}^{-}_{n}\right\}$, the central extention term
$i\oint_{S}\bm{k}_{\xi^{-}_{m}}(\delta_{\xi^{-}_{n}}\bar{\Phi},\bar{\Phi})=\frac{1}{12}\frac{3\sqrt{1+s}l}{G}m^{3}\delta_{m+n,0}$ (51)
where $\tilde{L}^{+}_{m}\equiv H_{\xi^{+}_{m}}$, $\tilde{L}^{-}_{m}\equiv H_{\xi^{-}_{m}}$, and $\bar{\Phi}$ is the background (V.1). To contain centerless subalgebra for $m=-1,0,+1$, we shift the zero modes of the charges
$L^{+}_{m}=$ $\tilde{L}^{+}_{m}+\delta_{m,0}\frac{3\sqrt{1+s}l}{24G},$
$L^{-}_{n}=$ $\tilde{L}^{-}_{n}+\delta_{n,0}\frac{3\sqrt{1+s}l}{24G}.$ (52)
Then, the Poisson bracket algebra of the charges are two copies of the Virasoro algebra
$i\left\{L^{+}_{m},L^{+}_{n}\right\}=$ $(m-n)L^{+}_{m+n}+\frac{c}{12}m(m^{2}-1)\delta_{m+n,0},$
$i\left\{L^{+}_{m},L^{-}_{n}\right\}=$ $0,$
$i\left\{L^{-}_{m},L^{-}_{n}\right\}=$ $(m-n)L^{-}_{m+n}+\frac{c}{12}m(m^{2}-1)\delta_{m+n,0}.$ (53)
The two central charges are the same, and with
$c=\frac{3\sqrt{1+s}l}{G}.$ (54)
By using the Cardy formula Cardy:1986ie ; deBoer:2010ac , and modified by the LV factor $\sqrt{1+s}$, the microscopic entropy of the dual CFT is
$S_{CFT}=2\pi(\sqrt{1+s})\sqrt{\frac{c}{6}(L^{-}_{0}-\frac{c}{24})}\pm 2\pi(\sqrt{1+s})\sqrt{\frac{c}{6}(L^{+}_{0}-\frac{c}{24})},$ (55)
The zero modes of $\tilde{L}^{+}_{m}$ and $\tilde{L}^{-}_{n}$ for the BTZ-like black hole (IV) are
$\tilde{L}^{+}_{0}=$ $H_{\xi^{+}_{0}}=\frac{\sqrt{1+s}l}{8G}\left(\mu-\frac{j}{l}\right),$
$\tilde{L}^{-}_{0}=$ $H_{\xi^{-}_{0}}=\frac{\sqrt{1+s}l}{8G}\left(\mu+\frac{j}{l}\right).$ (56)
Thus, we can obtain
$S_{CFT}=\frac{(1+s)\pi r_{{}_{\mathrm{H}}}}{G}.$ (57)
This result agrees precisely with the Bekenstein-Hawking entropy (41).
V.2 Warped conformal symmetries in near horizon region
In this subsection we will use the Detournay-Smoes-Wutte boundary conditions Detournay:2023zni to study the asymptotic symmetries of the near horizon extremal BTZ-like black hole.
In the extremal case, the BTZ-like metric and the bumblebee field (IV) are
$ds^{2}=$ $-\frac{(r^{2}-r^{2}_{h})^{2}}{l^{2}r^{2}}dt^{2}+\frac{(1+s)l^{2}r^{2}}{(r^{2}-r^{2}_{h})^{2}}dr^{2}+r^{2}(d\phi-\frac{r^{2}_{h}}{lr^{2}}dt)^{2},$
$B_{\mu}=$ $\left(0,~{}b_{0}\bigg{[}\frac{(1+s)l^{2}r^{2}}{(r^{2}-r^{2}_{h})^{2}}\bigg{]}^{\frac{1}{2}},~{}0\right),$ (58)
where $r_{h}=r_{+}=r_{-}=\frac{\sqrt{2}}{2}l\sqrt{\mu}$. We consider the change of coordinates
$t=\frac{\tau}{\varepsilon},\quad r^{2}=r^{2}_{h}+\varepsilon\rho,\quad\phi=\varphi+\frac{\tau}{\varepsilon l}.$ (59)
By taking limit $\varepsilon\rightarrow\infty$, we can obtain the near horizon geometry of the extremal BTZ-like metric and bumblebee field
$ds^{2}=$ $\frac{(1+s)l^{2}}{4\rho^{2}}d\rho^{2}+\frac{2\rho}{l}d\tau d\varphi+r^{2}_{h}d\varphi^{2},$
$B_{\mu}=$ $\left(0,~{}\frac{b_{0}\sqrt{1+s}l}{2\rho},~{}0\right).$ (60)
We now impose the Detournay-Smoes-Wutte boundary conditions Detournay:2023zni on the near horizon geometry of the extremal BTZ-like metric and bumblebee field (V.2), which is a higher dimensional uplift of Godet-Marteau boundary conditions in two-dimensional gravity Godet:2020xpk , and it is given by a finite coordinate transformation
$\tau\rightarrow\mathcal{F}(\tau),\quad\rho\rightarrow\frac{1}{\mathcal{F^{\prime}}}(\rho+\mathcal{G^{\prime}}(\tau)),\quad\varphi\rightarrow\varphi-\mathcal{G}(\tau).$ (61)
Here $\mathcal{F}(\tau)$ and $\mathcal{G}(\tau)$ are arbitrary functions of $\tau$. This transformation yields the metric and bumblebee field components
$g_{\tau\tau}=$ $r^{2}_{h}\mathcal{G^{\prime}}(\tau)^{2}-\frac{2}{l}\mathcal{G^{\prime}}(\tau)(\rho+\mathcal{G^{\prime}}(\tau))+\frac{(1+s)l^{2}(\mathcal{G^{\prime\prime}}(\tau)-\mathcal{H}(\tau)(\rho+\mathcal{G^{\prime}}(\tau)))^{2}}{4(\rho+\mathcal{G^{\prime}}(\tau))^{2}},$
$g_{\tau\rho}=$ $\frac{(1+s)l^{2}(\mathcal{G^{\prime\prime}}(\tau)-\mathcal{H}(\tau)(\rho+\mathcal{G^{\prime}}(\tau)))}{4(\rho+\mathcal{G^{\prime}}(\tau))^{2}},$
$g_{\tau\varphi}=$ $-r^{2}_{h}\mathcal{G^{\prime}}(\tau)+\frac{\rho+\mathcal{G^{\prime}}(\tau)}{l},$
$g_{\rho\rho}=$ $\frac{(1+s)l^{2}}{4(\rho+\mathcal{G^{\prime}}(\tau))^{2}},\quad g_{\rho\varphi}=0,\quad g_{\varphi\varphi}=r^{2}_{h},$
$B_{\tau}=$ $-\frac{b_{0}\sqrt{1+s}l(\rho\mathcal{H}(\tau)+\mathcal{H}(\tau)\mathcal{G^{\prime}}(\tau)-\mathcal{G^{\prime\prime}}(\tau))}{2(\rho+\mathcal{G^{\prime}}(\tau))},$
$B_{\rho}=$ $\frac{b_{0}\sqrt{1+s}l}{2(\rho+\mathcal{G^{\prime}}(\tau))},$
$B_{\varphi}=$ $0,$ (62)
where $\mathcal{H}(\tau)\equiv\mathcal{F^{\prime\prime}}(\tau)/\mathcal{F^{\prime}}(\tau)$. The near horizon geometry (V.2) is restored when $\mathcal{G^{\prime}}(\tau)=0=\mathcal{H}(\tau)$.
The asymptotic killing vectors generating the transformations (61) are given by
$\xi=\epsilon(\tau)\partial_{\tau}-(\rho\epsilon^{\prime}(\tau)+\zeta^{\prime}(\tau))\partial_{\rho}+\zeta(\tau)\partial_{\varphi},$ (63)
where $\epsilon(\tau)$ and $\zeta(\tau)$ are two arbitrary functions of $\tau$. By applying the Lie derivative on the metric and bumblebee field, the variations of $\mathcal{G^{\prime}}(\tau)$ and $\mathcal{H}(\tau)$ are given by
$\delta_{\xi}\mathcal{G^{\prime}}(\tau)=$ $\epsilon^{\prime}(\tau)\mathcal{G^{\prime}}(\tau)+\epsilon(\tau)\mathcal{G^{\prime\prime}}(\tau)-\zeta^{\prime}(\tau),$
$\delta_{\xi}\mathcal{H}(\tau)=$ $\epsilon^{\prime}(\tau)\mathcal{H}(\tau)+\epsilon^{\prime\prime}(\tau)+\epsilon(\tau)\mathcal{H^{\prime}}(\tau).$ (64)
To obtain the symmetry algebra, we assume that $\tau$ is periodic with period $L\in i\mathbb{R}$, where $i$ is the imaginary unit. The asymptotic killing vector (63) can be divided into modes
$l_{n}=$ $\xi\left(\epsilon(\tau)=\frac{L}{2\pi}e^{2\pi in\tau/L},~{}\zeta(\tau)=0\right),$
$j_{n}=$ $\xi\left(\epsilon(\tau)=0,~{}\zeta(\tau)=\frac{L}{2\pi i}e^{2\pi in\tau/L}\right),$ (65)
where $n\in\mathbb{Z}$. Under the Lie bracket, these modes satisfy the warped Witt algebra
$i\left[l_{m},l_{n}\right]=$ $(m-n)l_{m+n},$
$i\left[l_{m},j_{n}\right]=$ $-nj_{m+n},$
$i\left[j_{m},j_{n}\right]=$ $0$ (66)
To obtain finite and non-zero charges (II), we take $\partial{\Sigma}$ to be the circle S of constant $(r,\varphi)$ and take the limit $r\rightarrow\infty$. By taking variations of the metric and bumblebee field
$\delta g_{\mu\nu}=$ $\frac{\partial g_{\mu\nu}}{\partial\mathcal{G^{\prime}}(\tau)}\delta\mathcal{G^{\prime}}(\tau)+\frac{\partial g_{\mu\nu}}{\partial\mathcal{G^{\prime\prime}}(\tau)}\delta\mathcal{G^{\prime\prime}}(\tau)+\frac{\partial g_{\mu\nu}}{\partial\mathcal{H}(\tau)}\delta\mathcal{H}(\tau),$
$\delta B_{\mu}=$ $\frac{\partial B_{\mu}}{\partial\mathcal{G^{\prime}}(\tau)}\delta\mathcal{G^{\prime}}(\tau)+\frac{\partial B_{\mu}}{\partial\mathcal{G^{\prime\prime}}(\tau)}\delta\mathcal{G^{\prime\prime}}(\tau)+\frac{\partial B_{\mu}}{\partial\mathcal{H}(\tau)}\delta\mathcal{H}(\tau),$ (67)
the corresponding charge variation (II) associated with $\xi$ can be given by
[…]
which is integrable. Under the background (V.2), the charges are given by
[…]
The Poisson bracket algebra (II) of $\hat{L}_{n}$ and $\hat{J}_{n}$ is a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, the symmetry algebra of a WCFT,
$i\left\{\hat{L}_{m},\hat{L}_{n}\right\}=$ $(m-n)\hat{L}_{m+n}+\frac{\hat{c}}{12}m^{3}\delta_{m+n,0},$
$i\left\{\hat{L}_{m},\hat{J}_{n}\right\}=$ $-n\hat{J}_{m+n}-i\hat{\kappa}m^{2}\delta_{m+n,0},$
$i\left\{\hat{J}_{m},\hat{J}_{n}\right\}=$ $\frac{\hat{k}}{2}m\delta_{m+n,0},$ (70)
with central charges
$\hat{c}=\frac{3\sqrt{1+s}l}{G},\quad\quad\hat{\kappa}=0,\quad\quad\hat{k}=\frac{\sqrt{1+s}lL^{2}\mu}{4\pi^{2}G}.$ (71)
The $\hat{c}$ is the same as the Virasoro central charge $c$ of the AdS/CFT correspondence.
VI conclusions and remarks
In this paper, we employ the SPSM to investigate the conserved charges of a rotating BTZ-like black hole in Einstein-bumblebee theory of gravity. We explicitly computed the black hole mass, angular momentum and entropy, demonstrating that these quantities are influenced by LV parameter. Additionally, we derived the first law of black hole thermodynamics within this framework. As an asymptotical AdS spacetime, we studied the AdS/CFT correspondence and obtained the asymptotic charge algebra, which are two copies of the Virasoro algebra with non-trivial central charges. By applying the Cardy formula, we computed the microscopic entropy of the dual CFT, and found it is in exact agreement with the Bekenstein-Hawking entropy calculated by SPSM. Furthermore, in the near horizon region, we used the Detournay-Smoes-Wutte boundary conditions to study the near horizon geometry of extremal BTZ-like black hole. We obtained a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, which is the symmetry algebra of a WCFT.
Evidently, the conserved charges and central charges are affected by the LV parameter. We speculate that a similar modification arises in other bumblebee black hole solutions Casana:2017jkc ; Maluf:2020kgf ; Jha:2020pvk ; Filho:2022yrk ; Xu:2022frb ; Liu:2024axg , and also affect the thermodynamic properties. Notably, in the near horizon geometry, the central charge $\hat{c}$ of the Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra is found to be identical to the central charge $c$ of the Virasoro algebra in AdS/CFT correspondence. This result suggests a potential dual holographic description linking extreme black hole to WCFT, providing an alternative means to derive the black hole entropy by this duality.
Acknowledgements.
The author would like to thank Dr. Mingzhi Wang and Dr. Feiyi Liu for helpful discussions, and especially grateful to Prof. Xiang-Hua Zhai for carefully reading and improving this manuscript.